# Vision Transformers Don’t Need Trained Registers

### v4

@@ A.3 Attention maps after intervening upon register neurons
We present attention maps from all layers for several example images, applying our intervention on OpenCLIP (Figure 11) and DINOv2 (Figure 12 and Figure 13). Our results show that intervening upon register neurons produces clean attention maps free of the high-norm artifacts.
@@
Figure 11: Attention maps for OpenCLIP with test-time register. We show the mean CLS attention maps from all layers for several input images. Outliers appear in layer 6 for the original model but do not appear with test-time registers, producing clean, interpretable attention maps.
@@
Figure 12: Attention maps for DINOv2 with test-time register (Layers 0-11). We show the mean CLS attention maps from layers 0-11 for several input images. As we do not intervene in these layers, there is no difference in the attention maps.
@@
Figure 13: Attention maps for DINOv2 with test-time register (Layers 12-23). We show the mean CLS attention maps from layers 12-23 for several input images. Artifacts appear in uniform regions around layer 18 for the original model. With test-time registers, attention maps become clean and reveal the images’ main objects.
A.4 Additional activation maps of register neurons
@@
While our primary focus is on layer 6 of OpenCLIP (Section 3.1), we also include additional activation maps highlighting neurons that strongly activate at the top outlier location. As shown in Figure 14, the most active neurons from earlier layers produce activation maps that align closely with the output patch norms. This consistent alignment supports our hypothesis in Section 3.2 that a small number of sparsely activating neurons determine outlier locations even before they fully emerge. In OpenCLIP, we observe that top-activating neurons in earlier layers causally influence those in later layers. In particular, interventions on neurons in layer 5 or earlier lead to automatic changes in layer 6 activations (see Figure 5).
@@
A.5 Full experiment results on test-time registers in OpenCLIP
In Section 4, we moved outlier activations to an added, test-time register in OpenCLIP and observed that it removes outliers from the image patches. After moving the outliers to a single test-time register, we measured the test-time register’s output norm, the maximum norm of image patch outputs, and the highest last-layer CLS attention, both to the test-time register and any image patch. We now present the full results in Figure 15 and find a nearly identical distribution to Figure 5, indicating that the image patch outliers have been moved to the test-time register.
Figure 15: Intervening on activations of register neurons effectively shifts outliers to test-time registers. For all register neurons, we copy their highest activation into the test-time register and zero out the activations elsewhere. The test-time register absorbs the high norms and attention from the CLS token, indicating that the image patches no longer have outliers.
A.6 Additional experiments on register neurons in OpenCLIP
Intervening on random neurons is ineffective for shifting outliers. While register neurons are highly effective for shifting outliers (Section 3), we evaluate a random baseline to further justify our selection of neurons. We perform a similar intervention method used for register neurons on random neurons in OpenCLIP ViT-B/16, but we find they are ineffective for shifting outliers. For each layer, we randomly select the same number of neurons as those in our corresponding set of register neurons. During the forward pass, we intervene on the activation maps of these random neurons. We copy the highest activation across the original activation map to a randomly selected patch. In Figure 16, we observe that the selected patch does not absorb the high norms, demonstrating that using random neurons fails to shift outliers.
@@ A.6 Additional experiments on register neurons in OpenCLIP
Intervening on random neurons with high activations is similarly ineffective. We also test the hypothesis that it is the high activations specifically of register neurons that cause the outliers to appear. Instead of using the highest activation for the random neuron, we instead copy the highest activation from the corresponding register neuron. However, we find similarly ineffective results in Figure 17. The inability of random neurons to shift outliers reinforces our hypothesis that register neurons (i.e., their decoder direction) specifically are crucial for outlier emergence.
Figure 17: Copying in high activations into random neurons does not shift outliers. To verify that it is not the high activations of register neurons that create outliers, we intervene upon random neurons but copy in the highest activation from a corresponding register neuron. However, we find that the norm of the selected patch remains low, supporting our selection of register neurons to intervene upon.
Decoders of register neurons have large weights in certain dimensions. Sun et al. (2024) observe that outlier patches have massive activations in certain dimensions. Given this observation, we extract the decoder weights (post-non-linearity) for four register neurons from layer 5 and plot the distribution of weight magnitudes across dimensions in Figure 18. We find that certain dimensions (e.g. 579, 408) consistently have large weights across register neurons, which we do not observe with other random neurons. This property further supports the hypothesis that register neurons are the key mechanism that leads to outlier formation.
@@
A.7 Investigating outliers in DINOv2
We outline a similar investigation as presented in Section 3 on DINOv2-L/14 and find a sparse set of neurons that can be used to move outliers to arbitrary patches.
Outlier patches appear after MLPs. To evaluate whether the attention block or MLP causes outliers to form, we plot the maximum norm of the residual stream across all patches for every attention and MLP component. In Figure 19, we observe that the outliers appear at the MLP of layer 17, with attention sinks emerging soon after. These observations mirror Figure 2, where we also find a single layer in OpenCLIP that outliers tend to start appearing.
Figure 19: Outlier patches appear after MLPs in DINOv2; attention sinks appear after outlier patches. Left: Max norms across image patches (DINOv2-L/14). Right: max attention scores of the CLS token in the last layer. In both plots, we average across 1000 images. The increase in max norms and emergence of attention sinks occur in consecutive layers.
A small subset of neurons shows consistently high activations before outlier patches. To investigate layer 17, we evaluate the distribution of activations for its MLP neurons on the top outlier patch, averaged across 1000 images. To identify the top outlier patch, we find the patch with the maximum norm in the 2nd-to-last layer’s output (layer 22). We choose the 2nd-to-last layer to identify outliers because the maximum patch norm drops in the last layer output (Figure 19), making it difficult to differentiate outliers from normal patches. In Figure 20, we observe that the activation distribution for outliers is heavily skewed, with <25 neurons having disproportionately high activations. We find a similar skew in OpenCLIP (Figure 3), suggesting that outliers form due to a sparse set of neurons across different ViTs.
@@ A.7 Investigating outliers in DINOv2
Highly activated neurons activate on all outlier locations. Having seen that a small set of neurons highly activate on the top outlier, we now investigate whether these neurons highly activate across all outliers. We show three qualitative examples of activation maps in Figure 21 that closely align with the outliers. This alignment suggests that these highly activating neurons are responsible for outliers generally, which corroborates the results from OpenCLIP (Figure 4).
@@
Figure 21: Highly activated neurons on the top outlier in DINOv2 activate on all outlier positions. We present activation maps of three neurons from layer 17 that activate highly on the top outlier patch (Figure 3). These maps near-perfectly align with the high-norm outliers ("patch norms").
@@
Detecting register neurons. Given that these highly activating neurons–which we refer to as “register neurons”–appear responsible for outlier formation, we develop an automatic discovery algorithm to identify and intervene upon them to shift outliers. To identify register neurons (Algorithm 1), we compute the average activations at outlier patches for all MLP neurons and return the top neurons. To move outliers to arbitrary patches, we follow the same intervention method from Section 3.2. For each register neuron, we copy the highest activation across patches into the selected patch and zero out the activations elsewhere. For applying Algorithm 1 to DINOv2, we set $top\_k=45$, $highest\_layer=17$, and the outlier threshold to 150.
@@
Register neurons causally set the position of outliers. Following Section 3.2, we evaluate whether our intervention method can move outliers to arbitrary patches in DINOv2. After intervening, we measure the max norm of the selected patch, max norm of all other patches, and the highest last-layer CLS attentions to the selected patch and any other patch. We show in Figure 22 that our intervention successfully causes any selected patch to absorb the high norms and attention sinks. In Figure 1, we use register neurons to shift attention artifacts to form arbitrary spatial patterns, further demonstrating our ability to control outliers. These findings mirror that of OpenCLIP (Figure 5), exhibiting the generalizability of our intervention across ViTs.
@@
Figure 22: Moving outliers in DINOv2 to random patches. For all register neurons, we copy their highest activation into a selected patch and zero out the activations elsewhere. Left: norm of chosen random patch (yellow) and max norm of other patches (blue). Right: CLS attention to chosen random patch (yellow) and max CLS attention (blue) to any other patch. We find that the outliers shift to randomly selected patches, similar to OpenCLIP (Figure 5).
A.8 Extended Dense Prediction Results
We present additional dense prediction results from the experiments in Section 5.1. In Table 11, we show mean intersection over union (IOU), mean accuracy of each class (mAcc), and all pixel accuracy (aAcc) for ADE20k segmentation. In Table 11, we provide additional metrics for monocular depth estimation on the NYUv2 dataset. The rmse score measures root mean-squared error, while AbsRel is the absolute difference between the predicted depth and the true depth, relative to the true depth itself. $\delta_{1}$, $\delta_{2}$, and $\delta_{3}$ accuracy are defined as the percentage of pixels where the predicted depth is within $1.25$, $1.25^{2}$, and $1.25^{3}$ times the true depth, respectively. Models with test-time registers maintain or slightly improve performance across all metrics, with comparable performance to models trained explicitly with registers.
mIoU mAcc aAcc
DINOv2 ViT-L/14 48.2 59.3 81.9
w/ trained registers 49.1 60.8 81.7
w/ test-time register 49.1 61.5 81.9
OpenCLIP ViT-B/16 40.2 52.3 76.1
w/ test-time register 40.3 52.7 76.4
Table 10: Linear probe results for semantic segmentation on the ADE 20k dataset.
rmse $\downarrow$ AbsRel $\downarrow$ $\delta_{1}\uparrow$ $\delta_{2}\uparrow$ $\delta_{3}\uparrow$
DINOv2 ViT-L/14 38.7 11.3 89.0 98.9 99.8
w/ trained registers 38.2 11.0 89.1 98.9 99.8
w/ test-time register 37.8 11.0 89.5 98.9 99.8
OpenCLIP ViT-B/16 60.3 18.9 70.1 93.4 98.6
w/ test-time register 59.5 18.5 70.7 93.7 98.7
Table 11: Linear probe results for monocular depth estimation on the NYUv2 dataset.
A.9 Additional LOST object discovery results
@@
Visualizing Intermediate LOST Steps. We visualize the impact of test-time registers on the intermediate computation steps of the LOST object discovery algorithm in Figure 23. The first row (LOST score) displays the inverse degree of each patch in the similarity graph $\mathcal{G}$. The second row shows the similarity between all patch features and the initial seed. The third row highlights the initial seed in yellow and illustrates the resulting seed expansion. For DINOv2, adding a test-time register cleans up the intermediate maps used during LOST, translating to the performance gains in Table 5. Notably, the resulting intermediate steps resemble those obtained when using trained registers. Adding a test-time register to OpenCLIP also refines the intermediate maps. However, since OpenCLIP already produces higher-quality intermediate LOST maps compared to DINOv2, the resulting improvement in unsupervised object discovery is marginal, which was also observed by Darcet et al. (2024).
OpenCLIP LOST Feature Analysis. We visualize the LOST score for OpenCLIP using the key, query, and value projection features in Figure 24. Adding a test-time register results in maps that are more focused on the object. However, the value projection from the original model already filters out much of the noise in the background regions, making the baseline maps relatively clean. As a result, the improvement from using a test-time register is less pronounced compared to DINOv2. Darcet et al. (2024) suggest that for OpenCLIP, this may be the case since outliers seem to reside in the null space of the value projection layer.
@@
A.10 Additional VLM results
We present additional visualizations of LLaVA-Llama-3-8B patch norms and attention maps in Figure 25. As in Section 5.4, we visualize the patch norm map from the final layer of the vision encoder prior to projection into the language model input space. The patch norm maps highlight the presence of a sparse set of high-norm tokens, if any. We then visualize the average attention across all layers and heads of the language model from the response token to the visual tokens. We find that adding a test-time register to the vision encoder of the VLM leads to more interpretable attribution of the text output to the visual tokens.
Figure 25: Additional visualizations of LLaVA-Llama-3-8B patch norms and attention maps. As in Figure 7, we show patch norms of the vision encoder and the average attention from the answer token to the visual tokens. Adding a test-time register mitigates high-norm artifacts in the vision encoder which would otherwise lead to anomalous attention patterns in the language model.
A.11 Removing outliers with attention biases
We propose to create these attention biases in OpenCLIP ViT-B/16 training-free by approximating the effect of a test-time register on the attention computations. To set $\mathbf{k^{\prime}}$ and $\mathbf{v^{\prime}}$ for each attention head, we compute the mean key and value vectors of a test-time register token across 1000 images. At test-time, we zero out the activations of register neurons using the parameters from Section 3.2.
Results. We report zero-shot ImageNet performance in Table 12. Zeroing out the register neurons leads to a significant performance drop, but this drop is fully recovered by introducing the attention bias, which restores the model’s original performance. This suggests that zeroing out registers effectively removes outliers, while the attention bias approximates their influence at test time. Importantly, we no longer observe high-norm outliers across all layers (Figure 26). We also find that attention biases create clean attention maps free of artifacts comparable to using a test-time register (Figure 27). Our findings support the claim by Sun et al. (2024) that registers primarily function as attention biases. Although massive activations are relatively modest in vision transformers (e.g., < 500), they can exceed the mean activation by up to 10,000$\times$ in language models, creating challenges for quantization—where specialized methods are often needed to handle such outliers (Xiao et al. (2023)). Our results suggest that these large activations can be controlled without any additional training overhead. We leave the application of our method to the language domain for future work.
OpenCLIP Acc.
ViT-B/16 71.3
+ zeroed register neurons 55.6
+ attention bias 71.3
Figure 27: Using attention biases produces similarly clean attention maps as using a test-time register. We compute the attention map of the CLS token averaged across the attention heads of the last layer on OpenCLIP-B/16, finding that both attention biases and test-time registers remove artifacts.

### v5

@@ A.9 Attention maps after intervening upon register neurons
We present attention maps from all layers for several example images, applying our intervention on OpenCLIP (Figure˜16) and DINOv2 (Figure˜17 and Figure˜18). Our results show that intervening upon register neurons produces clean attention maps free of the high-norm artifacts.
@@
Figure 16: Attention maps for OpenCLIP with test-time register. We show the mean [CLS] attention maps from all layers for several input images. Outliers appear in layer 6 for the original model but do not appear with test-time registers, producing clean, interpretable attention maps.
@@
Figure 17: Attention maps for DINOv2 with test-time register (Layers 0-11). We show the mean [CLS] attention maps from layers 0-11 for several input images. As we do not intervene in these layers, there is no difference in the attention maps.
@@
Figure 18: Attention maps for DINOv2 with test-time register (Layers 12-23). We show the mean [CLS] attention maps from layers 12-23 for several input images. Artifacts appear in uniform regions around layer 18 for the original model. With test-time registers, attention maps become clean and reveal the images’ main objects.
A.10 Additional activation maps of register neurons
@@
While our primary focus is on layer 6 of OpenCLIP (Section˜3.1), we also include additional activation maps highlighting neurons that strongly activate at the top outlier location. As shown in Figure˜19, the most active neurons from earlier layers produce activation maps that align closely with the output patch norms. This consistent alignment supports our hypothesis in Section˜3.2 that a small number of sparsely activating neurons determine outlier locations even before they fully emerge. In OpenCLIP, we observe that top-activating neurons in earlier layers causally influence those in later layers. In particular, interventions on neurons in layer 5 or earlier lead to automatic changes in layer 6 activations (see Figure˜4).
@@
A.11 Additional OpenCLIP experiments
A.11.1 How do outliers emerge in ViTs?
A small subset of neurons have a consistently high contribution at the top outlier position. To localize the set of neurons that drive outlier formation, we evaluate the norm contribution of individual neurons at layer 6, where outliers appear. In Figure˜20, we plot the distribution of activations for all layer 6 neurons, comparing the average activation at a top outlier patch with a randomly selected non-outlier patch. We observe that five MLP neurons consistently have large activations on the top outlier, creating a skewed distribution, and low activations on non-outliers. To measure the effect of these neurons, we track their aggregate contribution to the residual stream over 1000 images by computing the norm and pairwise cosine similarities of their update vectors. As shown in Figure˜21, these neurons have consistent, high-norm contributions with effectively constant directions. In contrast, a random set of five neurons produces low-norm updates whose directions vary substantially.
Figure 20: Neuron activation distributions differ between outlier and non-outlier patches. For all layer 6 neurons, we average their activations on the top outlier patch and a randomly selected non-outlier patch across 1000 images. Non-outlier patches have a more symmetric distribution (left), whereas outlier patches show a skewed distribution with most activations near zero. Five neurons consistently exhibit high activations for outlier patches across images (right).
Figure 21: Outlier neurons show stable, high-magnitude contributions. We track the overall residual stream update from five random neurons and five identified outlier neurons on the top outlier patch across 1000 images. Outlier neurons contribute high norm updates (left) that are consistent in direction (right).
A.11.2 Adding register at test-time
In Section˜4, we moved outlier activations to an added, test-time register in OpenCLIP and observed that it removes outliers from the image patches. After moving the outliers to a single test-time register, we measured the test-time register’s output norm, the maximum norm of image patch outputs, and the highest last-layer [CLS] attention, both to the test-time register and any image patch. We now present the full results in Figure˜22 and find a nearly identical distribution to Figure˜4, indicating that the image patch outliers have been moved to the test-time register.
Figure 22: Intervening on activations of register neurons effectively shifts outliers to test-time registers. For all register neurons, we copy their highest activation into the test-time register and zero out the activations elsewhere. The test-time register absorbs the high norms and attention from the [CLS] token, indicating that the image patches no longer have outliers.
A.11.3 Intervening on random neurons
Intervening on random neurons is ineffective for shifting outliers. While register neurons are highly effective for shifting outliers (Section˜3), we evaluate a random baseline to further justify our selection of neurons. We perform a similar intervention method used for register neurons on random neurons in OpenCLIP ViT-B/16, but we find they are ineffective for shifting outliers. For each layer, we randomly select the same number of neurons as those in our corresponding set of register neurons. During the forward pass, we intervene on the activation maps of these random neurons. We copy the highest activation across the original activation map to a randomly selected patch. In Figure˜23, we observe that the selected patch does not absorb the high norms, demonstrating that using random neurons fails to shift outliers.
@@ A.11.3 Intervening on random neurons
Intervening on random neurons with high activations is similarly ineffective. We also test the hypothesis that it is the high activations specifically of register neurons that cause the outliers to appear. Instead of using the highest activation for the random neuron, we instead copy the highest activation from the corresponding register neuron. However, we find similarly ineffective results in Figure˜24. The inability of random neurons to shift outliers reinforces our hypothesis that register neurons (i.e., their decoder direction) specifically are crucial for outlier emergence.
Figure 24: Copying in high activations into random neurons does not shift outliers. To verify that it is the high activations in the corresponding register neuron channels that create outliers, we intervene upon random neurons but copy in the highest activation value from a corresponding register neuron. We find that the norm of the selected patch remains low, supporting our selection of register neurons to intervene upon.
Decoders of register neurons have large weights in certain dimensions. Sun et al. (2024) observe that outlier patches have massive activations in certain dimensions. Given this observation, we extract the decoder weights (post-non-linearity) for four register neurons from layer 5 and plot the distribution of weight magnitudes across dimensions in Figure˜25. We find that certain dimensions (e.g. 579, 408) consistently have large weights across register neurons, which we do not observe with other random neurons. This property further supports the hypothesis that register neurons are the key mechanism that leads to outlier formation.
@@
A.12 Investigating outliers in DINOv2
We outline a similar investigation as presented in Section˜3 on DINOv2-L/14 and find a sparse set of neurons that can be used to move outliers to arbitrary patches.
Outlier patches appear after MLPs. To evaluate whether the attention block or MLP causes outliers to form, we plot the maximum norm of the residual stream across all patches for every attention and MLP component. In Figure˜26, we observe that the outliers appear at the MLP of layer 17, with attention sinks emerging soon after. These observations mirror Figure˜2, where we also find a single layer in OpenCLIP that outliers tend to start appearing.
Figure 26: Outlier patches appear after MLPs in DINOv2; attention sinks appear after outlier patches. Left: Max norms across image patches (DINOv2-L/14). Right: max attention scores of the [CLS] token in the last layer. In both plots, we average across 1000 images. The increase in max norms and emergence of attention sinks occur in consecutive layers.
A small subset of neurons shows consistently high activations before outlier patches. To investigate layer 17, we evaluate the distribution of activations for its MLP neurons on the top outlier patch, averaged across 1000 images. To identify the top outlier patch, we find the patch with the maximum norm in the 2nd-to-last layer’s output (layer 22). We choose the 2nd-to-last layer to identify outliers because the maximum patch norm drops in the last layer output (Figure˜26), making it difficult to differentiate outliers from normal patches. In Figure˜27, we observe that the activation distribution for outliers is heavily skewed, with <25 neurons having disproportionately high activations. We find a similar skew in OpenCLIP (Figure˜20), suggesting that outliers form due to a sparse set of neurons across different ViTs.
@@ A.12 Investigating outliers in DINOv2
Highly activated neurons activate on all outlier locations. Having seen that a small set of neurons highly activate on the top outlier, we now investigate whether these neurons highly activate across all outliers. We show three qualitative examples of activation maps in Figure˜28 that closely align with the outliers. This alignment suggests that these highly activating neurons are responsible for outliers generally, which corroborates the results from OpenCLIP (Figure˜3).
@@
Figure 28: Highly activated neurons on the top outlier in DINOv2 activate on all outlier positions. We present activation maps of three neurons from layer 17 that activate highly on the top outlier patch. These maps near-perfectly align with the high-norm outliers ("patch norms").
@@
Detecting register neurons. Given that these highly activating neurons–which we refer to as “register neurons”–appear responsible for outlier formation, we develop an automatic discovery algorithm to identify and intervene upon them to shift outliers. To identify register neurons (Algorithm˜1), we compute the average activations at outlier patches for all MLP neurons and return the top neurons. To move outliers to arbitrary patches, we follow the same intervention method from Section˜3.2. For each register neuron, we copy the highest activation across patches into the selected patch and zero out the activations elsewhere. For applying Algorithm˜1 to DINOv2, we set top_k=45top\_k=45, highest_layer=17highest\_layer=17, and the outlier threshold to 150.
@@
Register neurons causally set the position of outliers. Following Section˜3.2, we evaluate whether our intervention method can move outliers to arbitrary patches in DINOv2. After intervening, we measure the max norm of the selected patch, max norm of all other patches, and the highest last-layer [CLS] attentions to the selected patch and any other patch. We show in Figure˜29 that our intervention successfully causes any selected patch to absorb the high norms and attention sinks. In Figure˜1, we use register neurons to shift attention artifacts to form arbitrary spatial patterns, further demonstrating our ability to control outliers. These findings mirror that of OpenCLIP (Figure˜4), exhibiting the generalizability of our intervention across ViTs.
@@
Figure 29: Moving outliers in DINOv2 to random patches. For all register neurons, we copy their highest activation into a selected patch and zero out the activations elsewhere. Left: norm of chosen random patch (yellow) and max norm of other patches (blue). Right: [CLS] attention to chosen random patch (yellow) and max [CLS] attention (blue) to any other patch. We find that the outliers shift to randomly selected patches, similar to OpenCLIP (Figure˜4).
A.13 Extended Linear Probe Results
We present detailed linear probe results in Table˜13 based on the experiments from Section˜5.1. We run 3 independent seeds for our linear probe evaluations and report 95% confidence intervals below. Performance differences remain consistent. Models with test-time registers maintain or slightly improve performance across all metrics, with comparable performance to models trained explicitly with registers.
IN ADE20k NYUd
Top-1 $\uparrow$ mIoU $\uparrow$ rmse $\downarrow$
DINOv2 ViT-L/14 86.36 ± 0.11 48.27 ± 0.15 0.3876 ± 0.0014
w/ trained registers 86.69 ± 0.09 49.07 ± 0.12 0.3823 ± 0.0011
w/ test-time register 86.38 ± 0.10 49.13 ± 0.11 0.3774 ± 0.0013
OpenCLIP ViT-B/16 77.42 ± 0.08 40.14 ± 0.11 0.6025 ± 0.0017
w/ test-time register 77.53 ± 0.07 40.29 ± 0.09 0.5956 ± 0.0016
A.14 Additional LOST object discovery results
@@
Visualizing Intermediate LOST Steps. We visualize the impact of test-time registers on the intermediate computation steps of the LOST object discovery algorithm in Figure˜30. The first row (LOST score) displays the inverse degree of each patch in the similarity graph $\mathcal{G}$. The second row shows the similarity between all patch features and the initial seed. The third row highlights the initial seed in yellow and illustrates the resulting seed expansion. For DINOv2, adding a test-time register cleans up the intermediate maps used during LOST, translating to the performance gains in Table˜5. Notably, the resulting intermediate steps resemble those obtained when using trained registers. Adding a test-time register to OpenCLIP also refines the intermediate maps. However, since OpenCLIP already produces higher-quality intermediate LOST maps compared to DINOv2, the resulting improvement in unsupervised object discovery is marginal, which was also observed by Darcet et al. (2024).
OpenCLIP LOST Feature Analysis. We visualize the LOST score for OpenCLIP using the key, query, and value projection features in Figure˜31. Adding a test-time register results in maps that are more focused on the object. However, the value projection from the original model already filters out much of the noise in the background regions, making the baseline maps relatively clean. As a result, the improvement from using a test-time register is less pronounced compared to DINOv2. Darcet et al. (2024) suggest that for OpenCLIP, this may be the case since outliers seem to reside in the null space of the value projection layer.
@@
A.15 Additional VLM results
We present additional visualizations of LLaVA-Llama-3-8B patch norms and attention maps in Figure˜32. As in Section˜5.4, we visualize the patch norm map from the final layer of the vision encoder prior to projection into the language model input space. The patch norm maps highlight the presence of a sparse set of high-norm tokens, if any. We then visualize the average attention across all layers and heads of the language model from the response token to the visual tokens. We find that adding a test-time register to the vision encoder of the VLM leads to more interpretable attribution of the text output to the visual tokens.
Figure 32: Additional visualizations of LLaVA-Llama-3-8B patch norms and attention maps. As in Figure˜6, we show patch norms of the vision encoder and the average attention from the answer token to the visual tokens. Adding a test-time register mitigates high-norm artifacts in the vision encoder which would otherwise lead to anomalous attention patterns in the language model.